OpenStudy (anonymous):

how do i change .00000000067 into a scientific notation?

OpenStudy (amistre64):

if you work with the people trying to help you, you might be able to catch on to these instead of making us feel like we are doing your homework/test for you.

OpenStudy (anonymous):

Again, as I said, there should be only one digit before decimal point. So you would write it as 6.7. Now for the exponent. The base should always be 10. 6.7 * 10^-10(because this is a decimal)

Parth (parthkohli):

0.00000000067 in a fraction is: \(\Large \color{MidnightBlue}{\Rightarrow {67 \over 100000000000 }\) Now, you can say, that it'd be: \(\Large \color{MidnightBlue}{\Rightarrow 67 * 10^-11 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6.7 * 10^-10 }\)

OpenStudy (lgbasallote):

hah fail \(\LaTeX\)

Parth (parthkohli):

Sorry :/

Parth (parthkohli):

It wasn't showing the preview

OpenStudy (lgbasallote):

there shouldnt be { before \large

Parth (parthkohli):

You basically shift the places. Got it now?

OpenStudy (anonymous):

thanks for those who helped instead of judging! i havent been in school in 8 years sooo bare with me, smartasss!! (amistre64)

Parth (parthkohli):

For more help, see this http://www.youtube.com/watch?v=i6lfVUp5RW8

OpenStudy (amistre64):

1 hour suspension

OpenStudy (apoorvk):

Significant digits ---> the no. of digits after the decimal is known as the no. of significant digits. so, when you shift the decimal by suppose 'n' places to the right (or right), you multiply the no. by "10^(-n) (or 10^-n) if moving to the left)to maintain equivalence. And in scientific notation, the convention is such that we write ONE significant digit before the decimal, the rest after it, and then for the decimals, multiply an appropriate "10^(-n)" over there. so, in your case your no. changes to: \[67 \times 10^{-11}\] \[=6.7 \times 10^{-10}\] As another example, \[ 34,000 = 3.4 \times 10^4\] helps?